OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)?

OpenStudy (anonymous):

m= slope = \[\frac{y2-y1}{x2-x1}\] \[m=\frac{7-(-8)}{3-(-2)} =5\] Parralel lines have the same slope., so the slope of both lines is 5

OpenStudy (anonymous):

If you need to write an equation for the parallel line: \[y-y1=m(x-x1)\] \[y-3=5(x+2)\] \[y=5x+10+3=5x+13\]